Find the integral surface of the partial differential equation:(x–y) y^2 p+(y–x)x^2 q=(x^2 +y^2)z through the curve xz=a^2, y=0
1
Expert's answer
2020-03-02T09:00:58-0500

"(x-y)y^2p + (y-x)x^2q=(x^2+y^2)z" is the equation to be solved for the curves

"xz=a^2" ---(1)

"y=0" ---(2)

Lagrange's auxillary equation is :

"dx\/((x-y)y^2) = dy\/((y-x)x^2)=dz\/((x^2+y^2)z)"

Solving the first two sides; we get;

"\int (dx/x^2)= -\int (dy/y^2)"

"\implies x^3+y^3=c_1" ---(3)


Now using "-1\/(x-y),1\/(x-y);1\/z" as Lagrange multipliers, we get;

"-dx\/(x-y)+dy\/(x-y)+dz\/z=0"

"\implies dz/z=dx-dy/(x-y)"

Let "x-y=t \implies dx-dy=dt"

"\therefore \int dz/z= \int dt/t"

On solving we get;

"t/z=c_2 \implies (x-y)/z=c_2" ---(4)


Using (2) in (3) and (4) we get;

"x_3=c_1" ---(5)

"x=c_2z"

Now using (1) to substitute for z in the above equation, we get;

"x=ac^{1\/2}" ---(6)

Using (6) in (5), we get;

"c_1^2=c_2^3a^6" ---(7)

Now substituting "c_1, c_2" in (7) from (3) and (4), we get;

"(x^3+y^3)^2=(x-y)^3a^6\/z^3" ----(Answer)